(D2-D')z=ex-ysin(x+2y)


1
Expert's answer
2021-06-13T16:02:13-0400

Given differential equation is "(D^2-D')z = e^{x-y}sin(x+2y)"

Let "D' = 1"

Then auxiliary equation will be, "m^2-1=0 \implies m = 1,-1"


CF of the equation, "z = f_1(y+x)+f_2(y-x)"


"PI = \frac{1}{D^2-D'}e^{x-y}sin(x+2y)" "= e^{x-y}\frac{1}{(D+1)^2-(D'-1)}sin(x+2y)"

"= e^{x-y}\frac{1}{D^2+2D-D'+2}sin(x+2y)" "= e^{x-y}\frac{1}{-1+2D-D'+2}sin(x+2y)"


"= e^{x-y}\frac{1}{2D-D'+1}sin(x+2y)"


Now Multiplying by D in numerator and denominator,


"PI= e^{x-y}\frac{D}{2D^2-D'D+D}sin(x+2y)"


Putting values, "D^2 = -1, DD' = -2"


"PI= e^{x-y}\frac{D}{-2+2+D}sin(x+2y) = e^{x-y}sin(x+2y)"


Hence, complete solution will be,

"z = f_1(y+x)+f_2(y-x) + e^{x-y}sin(x+2y)"